OpenStudy (anonymous):

f(x)-x-8, find f(-10) so f(-10) = -10-8 =-18, is this correct?

OpenStudy (anonymous):

Well, did you mean f(x)=x-8 or f(x) = -x-8

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

f(x) = -(-10) -8 OR in other words, f(x) = -1(-10) - 8

OpenStudy (anonymous):

Can you solve for f(-10) now?

OpenStudy (anonymous):

so the solution is 2

OpenStudy (anonymous):

Correct. :) f(-10) = 2

OpenStudy (anonymous):

whoohoo! im a roll tonight, thank you so much jiji501 :)